Which of the following statements are true and which are false? Give reasons for your
answers, in the form of a short proof or a counter example.
(i) The function f defined by f(x) = tan(2x) is a periodic function with period Ο€.

(ii) The function f: R---> R defined by f(x) = 1-|x| is differentiable at x=1

(iii) The function f: [3,4] ----> R defined by
f(x) = x
1
Expert's answer
2021-09-29T12:43:03-0400

(i) Function f(x) = tanx is a periodic function with period "\pi".

So, period of tan2x is "\pi"/2. This is because period of tanax is "\pi"/a.


The given statement is false.


(ii) Lets graph f(x) = 1 - |x|



Clearly, at x = 1, there is only tangent. So, the function is differentiable at x = 1.


The given statement is true.


(iii) Lets graph the function f(x) = x





Clearly, the function is defined in the given domain [3, 4].


So the given statement is correct.


The final graph of the function is given as,